OpenStudy (anonymous):

calculate the slope of the line passing through the given points. If the slope is undefined, so state. Then indicate whether the line rises, falls, is horizontal, or is vertical. (4,2) and (3,4)

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2) so when given the points (4,2) and (3,4) \[\frac{ 4-2 }{ 3-4 }\] simplify and that's your slope

OpenStudy (anonymous):

how do i find out if the line rises, falls, is horizontal, or is vertical ?

jigglypuff314 (jigglypuff314):

vertical is when you try to find the slope and you get a 0 in the denominator horizontal is when you get a 0 in the numerator it falls when you get a negative number it rises when you get a positive number

OpenStudy (anonymous):

how do you know when the slope is undefined ?